# Understanding differences in applying DETR to natural and medical images

### v1

@@ 5 Conclusions
In this study, we investigated the impact of common design choices in Deformable DETR on medical images object detection, focusing on the NYU Breast Cancer Screening Dataset. We found out that all the design choices we experimented with need to be reconsidered and a simpler architectures typically lead to better performance on medical dataset. Our results further conclude that while the model effectively identifies abnormal tissues, it struggles to distinguish benign from malignant lesions and assigning high classification scores to the latter. This issue partly stems from the relatively small proportion of positive cases, as only about $10\%$ of the images contain positive objects. Partly because visual features typically used for classification in natural image datasets, such as object size, shape, and color, are less discriminative in medical images. For example malignant lesions in our images often lack distinct borders, have uniform sizes, and are presented in grayscale.
Future research should prioritize developing techniques for extracting subtle features relevant to medical imaging, including texture variations, density changes, irregular borders, and microcalcifications. Additionally, exploring architectural designs capable of efficiently processing full-resolution images is crucial, enabling models to focus on rich information in the pertinent regions while disregarding extensive background areas. Additionally, implementing regularization methods to mitigate overfitting in classification tasks, especially in datasets with limited positive cases, is vital for improving model performance.
@@ 6 Limitations and future work
Our study had several limitations. First, while we focused on Deformable DETR, comparing its performance with other transformer-based object detection architectures, such as DINO, would provide a broader understanding of the strengths and weaknesses of different models. Future studies should explore multiple architectures to reach more generalizable conclusions. Additionally, due to the amount of the resources necessary to conduct the experiments with a large medical dataset, our investigation was confined to a single dataset. To ensure the robustness of our findings, it is important to validate these results across various medical imaging datasets. Considering the vast diversity within medical image datasets, including modalities such as brain MRI or ultrasound images, future research should encompass these distinct characteristics to provide a comprehensive evaluation.
Acknowledgments
This work was supported in part by a grants from the National Institutes of Health (P41EB017183), the National Science Foundation (1922658), the Gordon and Betty Moore Foundation (9683), and the Mary Kay Ash Foundation (05-22). We also gratefully acknowledge the support of Nvidia Corporation with the donation of some of the GPUs used in this research.
Ethical Standards
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
Conflicts of Interest
@@ A.1 Multi-head self-attention (MHSA)
$MHSA(Q,K,V)=\sum_{m=1}^{M}W_{mo}\left[\mathrm{softmax}\left(\frac{QW_{mq}(KW_{mk})^{T}}{\sqrt{d/M}}\right)VW_{mv}\right].$ (3)
@@
$W_{mq},W_{mk},W_{mv}\in\mathbb{R}^{d\times d/M}$ linearly transforms $K,Q,V$ in $m$-th head and $W_{mo}\in\mathbb{R}^{d/M\times d}$.
@@ A.2 Multi-head (MH) cross-attention
The MH cross-attention module performs the same computation as MHSA defined in 3, except that $K,Q,V$ are defined based on two different set of tokens. The queries $Q$ is defined by object queries $q=q_{c}+q_{p}$ where $q_{p}$ and $q_{c}$ are the positional embedding and content embedding of the object queries. The keys $K$ are defined by encoder features $x_{enc}+x_{p}$. Specifically,
@@ A.3 Set prediction loss
DETR uses a set prediction loss which enables end-to-end training without non-maximum suppression (NMS). DETR produces a fixed number of predictions per image $N$, and $N$ is set to be significantly larger than the maximum possible number of objects in the image. Let $\{\hat{y}_{i}=(\hat{c}_{i},\hat{b}_{i})\}_{N}$ be all pairs of class and box predictions. The set of $N$ labels is $\{y_{i}=(c_{i},b_{i})\}_{N}$ where each ground truth label represents an object in the image. If there are fewer objects than $N$, the rest of the labels are the empty classes $(0,\emptyset)$. The set prediction loss is computed in two steps. The first step is to find a permutation $\sigma$ on the set of labels $\{y_{i}\}$ that minimizes the matching loss, defined as below:
@@
$\mathcal{L}_{match}(\hat{y}_{i},y_{\sigma(i)})=W_{cls}\mathcal{L}_{cls}(\hat{c}_{i},c_{\sigma(i)})+1_{b\neq\emptyset}\left(W_{l1}\mathcal{L}_{l1}(\hat{b}_{i},b_{\sigma(i)})+W_{giou}\mathcal{L}_{giou}(\hat{b}_{i},b_{\sigma(i)})\right),$
where $W_{cls},W_{l1},W_{giou}$ are scalar coefficients that are tuned as hyperparameters to balance the scale of different losses. The Hungarian algorithm (Kuhn, 1955) can efficiently find the optimal matching $\hat{\sigma}$. The second step is to minimize the loss function $\sum_{N}^{i}\mathcal{L}_{match}(\hat{y}_{i},y_{\hat{\sigma}(i)})$ with the permutation $\hat{\sigma}$ on the label set.
DETR also utilizes auxiliary loss in each decoder layer to provide stronger supervision. At the end of each decoder layer, it predicts $N$ boxes and class scores with MLP prediction heads. All the prediction heads share weights. The above two steps, the matching step and Hungarian loss minimization, are applied to each decoder layer’s output. At inference, only the last layer’s output are used as the final prediction.
B Deformable DETR architecture
@@ B.1 Deformable multi-head self-attention
$\mathrm{Deformable\_MHSA}(Q_{q},K,V)=\sum_{m=1}^{M}W_{mo}\left[\mathrm{softmax}\left(K_{q}W_{mk}\right)V_{q}W_{mv}\right]$ (6)
@@
$Q=x_{f}+x_{p}\text{ and }Q_{q}=q,K_{q}=\delta(K,q)\in\mathbb{R}^{k\times d},V_{q}=\delta(V,q)\in\mathbb{R}^{k\times d}.$ (7)
The key sampling function $\delta$ samples $k$ keys from the full set of keys $K=x_{f}+x_{p}$ by generating the sampling offsets $\Delta p$ with respect to reference points $p_{q}$: $\delta(K,q)=K(p_{q}+\Delta p)$. The sampling offsets are obtained through a linear transformation of the query $q$.
C Iterative Bounding Box Refinement Technique
@@
Throughout the decoder, the locations of these reference points remain constant. They are updated based on the learnable positional embedding $p_{q}$ when a backward pass is completed. Formally, let $r_{q}^{i}$ be the reference points of an object query $q$ in i-th decoder layer. In standard Deformable DETR,
@@
In IBBR, the reference points $r_{q}^{i}$ in i-th decoder layer is refined based on the previous reference points $r_{q}^{i-1}$ and the offsets predicted by auxiliary prediction head.
@@
D Backbone Pre-training
We pretrained the Swin-T Transfromer backbone with a cancer classification task on our dataset. This classification task is a binary multi-label classification that predicts two scores indicating if an input image contains benign lesions or malignant lesions.
E Hyperparameter Tuning
Our method for hyper-parameter tuning is random search. We tuned the following hyperparameters and their ranges on quarter resolution images:
@@ E Hyperparameter Tuning
We train $80$ jobs in total and choose the best model based on FAUC_1.

### v2

@@ 5 Conclusions
In this study, we investigated the impact of common design choices in Deformable DETR (Zhu et al., 2020) on object detection performance in medical imaging, using two representative datasets: the NYU Breast Cancer Screening Dataset and LUNA16. We found that all the design choices we experimented with need to be reconsidered, and simpler architectures typically lead to better performance on medical dataset.
Additionally, our findings suggest that the model tends to struggle more with correctly classifying detected objects than with localizing them. Many design choices developed for natural image detection, such as query initialization, multi-scale feature fusion, and bounding box refinement, are primarily aimed at improving localization. However, since classification appears to be the more challenging component in medical imaging, these localization-focused techniques may offer limited benefit and, in some cases, even hinder performance.
Future research should focus on developing architectures specifically tailored to the characteristics of medical imaging. This includes improving the model’s ability to extract subtle visual cues that are often critical for classification, such as texture variations, tissue density changes, irregular borders, and microcalcifications. Another important direction is designing architectures that can efficiently process full-resolution images, allowing the model to leverage detailed information in relevant regions while minimizing the influence of background areas. Moreover, addressing overfitting in classification tasks, particularly in datasets with limited positive samples, requires the integration of effective regularization techniques to improve generalization and robustness.
@@ 6 Limitations and future work
Our study has several limitations. First, while our results demonstrate that simplified DETR configurations perform well on medical imaging tasks, future studies should explore additional architectural designs within the DETR family to validate and extend these findings, for example, constrastive denoising training in DINO (Zhang et al., 2022) and anchors in Anchor-DETR (Wang et al., 2022). Second, our experiments were conducted primarily on the NYU Breast Cancer Screening Dataset and LUNA16 for lung nodule detection. While these datasets capture important aspects of medical imaging, future studies should evaluate model performance across a broader set of imaging modalities and clinical tasks, such as brain MRI, ultrasound, or multi-phase CT, to assess the generalizability of our conclusions.
[…]
This work was supported in part by grants from the National Institutes of Health (P41EB017183), the National Science Foundation (1922658), the Gordon and Betty Moore Foundation (9683), and the Mary Kay Ash Foundation (05-22). We also appreciate the support of Nvidia Corporation with the donation of some of the GPUs used in this research.
[…]
This retrospective study was approved by the NYU Langone Health Institutional Review Board (ID#i18-00712_CR3) and is compliant with the Health Insurance Portability and Accountability Act. Informed consent was waived since the study presents no more than minimal risk.
[…]
@@ A.1 Multi-head self-attention (MHSA)
$\mathrm{MHSA}(Q,K,V)$ (3)
$=\sum_{m=1}^{M}W_{mo}\left[\mathrm{softmax}\left(\frac{QW_{mq}(KW_{mk})^{T}}{\sqrt{d/M}}\right)VW_{mv}\right].$ (4)
@@
$W_{mq},W_{mk},W_{mv}\in\mathbb{R}^{d\times d/M}$ linearly transforms $K,Q,V$ in the $m$-th head and $W_{mo}\in\mathbb{R}^{d/M\times d}$.
@@ A.2 Multi-head (MH) cross-attention
The MH cross-attention module performs the same computation as the MHSA defined in 3, except that $K,Q,V$ are defined based on two different sets of tokens. The queries $Q$ are defined by the object queries $q=q_{c}+q_{p}$, where $q_{p}$ and $q_{c}$ are the positional embedding and the content embedding of the object queries. The keys $K$ are defined by the encoder features $x_{enc}+x_{p}$. Specifically,
@@ A.3 Set prediction loss
DETR uses a set prediction loss that enables end-to-end training without non-maximum suppression (NMS). DETR produces a fixed number of predictions per image $N$, and $N$ is set to be significantly larger than the maximum possible number of objects in the image. Let $\{\hat{y}_{i}=(\hat{c}_{i},\hat{b}_{i})\}_{N}$ be all pairs of class and box predictions. The set of $N$ labels is $\{y_{i}=(c_{i},b_{i})\}_{N}$ where each ground truth label represents an object in the image. If there are fewer objects than $N$, the rest of the labels are empty classes $(0,\emptyset)$. The set prediction loss is computed in two steps. The first step is to find a permutation $\sigma$ on the set of labels $\{y_{i}\}$ that minimizes the matching loss, defined as below:
@@
$\mathcal{L}_{match}(\hat{y}_{i},y_{\sigma(i)})=W_{cls}\mathcal{L}_{cls}(\hat{c}_{i},c_{\sigma(i)})$
$+1_{b\neq\emptyset}\left(W_{l1}\mathcal{L}_{l1}(\hat{b}_{i},b_{\sigma(i)})+W_{giou}\mathcal{L}_{giou}(\hat{b}_{i},b_{\sigma(i)})\right),$
where $W_{cls},W_{l1},W_{giou}$ are scalar coefficients that are tuned as hyperparameters to balance the scale of different losses. The Hungarian algorithm (Kuhn, 1955) can efficiently find the optimal match $\hat{\sigma}$. The second step is to minimize the loss function $\sum_{N}^{i}\mathcal{L}_{match}(\hat{y}_{i},y_{\hat{\sigma}(i)})$ with the permutation $\hat{\sigma}$ on the label set.
DETR also utilizes auxiliary loss in each decoder layer to provide stronger supervision. At the end of each decoder layer, it predicts $N$ boxes and class scores with MLP prediction heads. All prediction heads share weights. The above two steps, the matching step and the Hungarian loss minimization, are applied to each decoder layer’s output. In inference, only the output of the last layer is used as the final prediction.
Appendix B Deformable DETR architecture
@@ B.1 Deformable multi-head self-attention
$\mathrm{Deform\_MHSA}(Q_{q},K,V)$ (7)
$=\sum_{m=1}^{M}W_{mo}\left[\mathrm{softmax}\left(K_{q}W_{mk}\right)V_{q}W_{mv}\right]$ (8)
@@
$Q=x_{f}+x_{p},$ (9)
$Q_{q}=q,$ (10)
$K_{q}=\delta(K,q)\in\mathbb{R}^{k\times d},$ (11)
$V_{q}=\delta(V,q)\in\mathbb{R}^{k\times d}.$ (12)
The key sampling function $\delta$ samples the $k$ keys from the full set of keys $K=x_{f}+x_{p}$ by generating the sampling offsets $\Delta p$ with respect to reference points $p_{q}$: $\delta(K,q)=K(p_{q}+\Delta p)$. The sampling offsets are obtained by linear transformation of the query $q$.
Appendix C Iterative Bounding Box Refinement Technique
@@
Throughout the decoder, the locations of these reference points remain constant. They are updated based on the learnable positional embedding $p_{q}$ when a backward pass is completed. Formally, let $r_{q}^{i}$ be the reference points of an object query $q$ in the i-th decoder layer. In standard Deformable DETR,
@@
In IBBR, the reference points $r_{q}^{i}$ in i-th decoder layer are refined based on the previous reference points $r_{q}^{i-1}$ and the offsets predicted by the auxiliary prediction head, which is a Multi-layer Perceptron (MLP). MLP is defined by three fully connected layers, which transform the output embeddings of the transformer into the desired bounding box coordinates.
@@
Appendix D Backbone Pre-training
We pretrained the Swin-T Transfromer backbone with a cancer classification task on our dataset. This classification task is a binary multi-label classification that predicts two scores indicating if an input image contains benign lesions and/or malignant lesions.
Appendix E Hyperparameter Tuning
Our method for hyper-parameter tuning is random search. We tuned the following hyperparameters and their ranges on quarter-resolution images:
@@ Appendix E Hyperparameter Tuning
We train $80$ jobs in total and choose the best model based on $\mathrm{FAUC}^{1}_{10}$.